### v1

# thanks: These two authors contributed equallythanks: These two authors contributed equally

Abstract
We carry out strain-controlled in situ compression experiments of micron-sized tungsten micropillars (W) in the temperature range 300-900K, together with simulations of three-dimensional discrete dislocation dynamics (DDD) on the same scale. Two distinct regimes are observed. At low temperatures, plastic deformation appears smooth, both temporally and spatially. Stress fluctuations are consistent with a Wiener stochastic process resulting from uncorrelated dislocation activity within the pillars. However, high-temperature stress fluctuations are highly correlated and exhibit features of self-organized criticality (SOC), with deformation located within well-defined slip bands. The high-temperature stress relaxation statistics are consistent with a thermally activated nucleation process from the surface. It is argued that the nature of the transition between the two regimes is the same as in the brittle to ductile transition in bulk BCC metals.
During the last two decades, acoustic emission and micropillar compression experiments have revealed that plastic deformation in crystalline materials manifests at the microscale in the form of intermittent dislocation activity (Weiss et al., 2001; Dimiduk et al., 2006; Ispánovity et al., 2022; Papanikolaou et al., 2017; Weiss and Marsan, 2003; Brinckmann et al., 2008; Zaiser et al., 2008). This observation suggests the intriguing hypothesis that, at a fundamental level, plasticity could be governed by the principle of self-organized criticality (SOC) Bak et al. (1987), a hypothesis already advanced within the context of other physical phenomena, such as solar flares Crosby et al. (1993), earthquakes Sornette and Sornette (1989), snow avalanches, and sandpile slides Miguel et al. (2001); Maaß and Derlet (2018). Compared to these fields, where experimental parameters are often impossible to change, microscale plasticity provided a unique platform to test one of the fundamental assumptions of SOC, that the cutoff of the size distribution depends only on the size of the system and not on external tunable parameters Marković and Gros (2014). In fact, micropillar experiments appeared to defy this assumption, revealing that external parameters such as stress Friedman et al. (2012) and strain rate Maaß and Derlet (2018) affect the size distribution cutoff, thus giving rise to the alternative concept of tuned criticality. It was proposed that the SOC-like behavior of certain crystals, such as ice, could be the limit of a continuum spectrum, with a “mild to wild” transition controlled by a single ratio $R=L/\ell$ between the size of the system $L$ and an intrinsic length scale $\ell=\mu b/\tau$ determined by the overall resistance $\tau$ to dislocation motion Weiss (2019). For example, the introduction of precipitates of high pinning strength in Al pillars was shown to inhibit strain bursts Zhang et al. (2017). More recently, micropillar compression tests carried out in body-centered cubic (BCC) crystals showed that the distribution of avalanche sizes either follows a truncated power law similar to face-centered cubic (FCC) crystals, but with a temperature-dependent cutoff Alcalá et al. (2020); Zhang et al. (2023), or exhibits an exponential to power-law transition as temperature increases Rizzardi et al. (2022). We posit that the role of temperature in relation to self-critical behavior in BCC metals remains unclear.
In this work, we show that low-temperature avalanches in W micropillars are consistent with a Wiener stochastic process resulting from uncorrelated dislocation activity within the pillars. In contrast, high-temperature stress fluctuations are highly correlated, exhibit characteristics of SOC, and are consistent with a different internal mechanism. We argue that the transition between these two regimes is a universal feature of BCC micropillars and that the nature of this transition is the same as in the brittle-to-ductile transition (BDT) of bulk BCC metals.
Figure 1: (a) Micropillar compressoin stress-strain curves from 300K to 900K. An incremental $1\%$ strain shift is applied to each temperature for clarity. The SEM images in the insert indicate the occurrence of a large strain burst. (b) Size of representative pre-compressed pillar and post-compression SEM images of micropillars at 300K, 600K and 800K.
[…]
Representative stress-strain curves are presented in Fig. 1, together with selected pre- and post-compression SEM images. The room temperature Young modulus $E\approx 260$ GPa and the strength $\sigma_{y}\approx 1.56$ GPa are consistent with previous measurements in W pillars of similar size Schneider et al. (2009); Abad et al. (2016). The diffuse slip is observed at low temperatures, while the localized $a/2\langle 111\rangle\{110\}$ slip bands become more pronounced as the temperature increases. The overall decrease in strength with increasing temperature is a signature of the thermally activated mobility of screw dislocations in BCC crystals. The striking feature of these stress-strain curves is the transition from a smooth plastic flow to abrupt stress drops with increasing temperature. This transition can be characterized by the distribution of the size of the dislocation avalanche. To construct avalanche statistics, we recall that the primary indicator of the dynamics of collective displacement is the plastic strain rate $\dot{\varepsilon}^{p}$, which is proportional to the stress rate $\dot{\sigma}\approx-E\dot{\varepsilon}^{p}$ when avalanches occur under a sufficiently slow displacement control. Following Laurson and Alava (2006); Papanikolaou et al. (2012), Wiener filtering was applied to the nanoindenter load cell signal, which, upon appropriate conversion to an engineering stress rate $\dot{\sigma}$, was filtered to extract avalanche events of duration $T$ and size $S=\int_{0}^{T}\dot{\sigma}\text{d}t$ (details can be found in the supplementary materials). Fig. 2a shows the complementary cumulative distribution functions (CCDF) of the event sizes for different temperatures. A truncated power law scaling is found to capture data above 600K, with an exponent of $\tau\sim 0.74$ for the burst size. This exponent agrees well with the value observed in many pure metals (between 0.5-0.9), including Ni, Au, Al, Cu, Mo, and Nb Dimiduk et al. (2006); Brinckmann et al. (2008); Zhang et al. (2017); Papanikolaou et al. (2017), as well as with the mean-field universality class Fisher (1998). With decreasing temperature, the size of events decreases drastically, and distributions spanning less than a decade are hardly attributed to critical behavior Stumpf and Porter (2012). A similar trend in the temperature dependence of the avalanche size distribution was also observed in other BCC micropillars such as Nb and Mo Rizzardi et al. (2022); Zhang et al. (2023).
Figure 2: (a) Complementary cumulative distribution function (CCDF) of avalanche size at different temperatures. The inset shows that the power law distribution of the “size” $\mathcal{S}$ agrees with a Wiener process at low temperature. (b) PSD of stress fluctuations at different temperatures. For clarity, a shift along the $y$axis is applied.
It should be recognized that power law distributions size are generated by many processes, including branching processes and random walks, and therefore they represent a necessary but not sufficient signature of critical phenomena Watkins et al. (2016); Jensen (2021). Due to the lack of dominant characteristic length and time scales, temporal (and spatial) correlations are the true hallmark of critical phenomena, manifesting as “crackling” $1/f$ noise in the power spectrum Kuntz and Sethna (2000); Laurson et al. (2005). Compared to the avalanche size analysis, the power spectral density (PSD) of stress fluctuations offers therefore more direct evidence of correlated dislocation activity within the micropillars. The PSD of the stress signal is shown in Fig. 2b. The PSD is well described by a power law $P_{\sigma}(f)\sim f^{-\gamma}$ at all temperatures, which implies $P_{\dot{\varepsilon}^{p}}(f)\sim f^{-\gamma+2}$ for the plastic strain rate. At temperatures above 600K, a power law exponent $\gamma\sim 3$ is observed, corresponding to a $\sim 1/f$ “crackling” noise in the plastic deformation rate. This indicates that high-temperature plasticity within the W micropillars is dominated by highly correlated dislocation activity. Consistent with the original intuition of SOC Bak et al. (1987), correlated activity manifests in heavy-tailed avalanche size distributions, as observed in Fig. 2a at high temperatures. By contrast, below 500K the stress PSD scales as a power law with exponent $\gamma\sim 2$, which is typical of a Wiener process (or a continuous Brownian motion). Note that the size CCDF of a Wiener process is a power law with exponent $1/3$ Villegas et al. (2019), which is in agreement with the distribution of “sizes” $\mathcal{S}=\int_{0}^{T}\sigma\text{d}t$ shown in the inset of Fig. 2a at low temperatures. A Wiener process represents additive white noise, that is $P_{\dot{\varepsilon}^{p}}\sim f^{0}$. Hence, low-temperature plasticity in the micropillars appears to be mediated by uncorrelated dislocation activity. This uncorrelated dislocation activity, which is an indication of negligible dislocation-dislocation interactions, also manifests itself in the lack of significant strain hardening.
Figure 3: (a) Stress-strain curves of DDD micropillar compression simulations at 300K and 900K using different initial microstructures. (b) Corresponding power spectrum of the flow stress from the simulations.
To elucidate the underlying mechanisms of low- and high-temperature plasticity in W micropillars, we carry out three-dimensional DDD simulations in cylindrical domains with the same dimensions and crystallographic orientations as the experimental samples. We employ the MoDELib package, which features a dislocation mobility law tuned to capture the thermally activated motion of screw dislocations in W Po et al. (2016); Cui et al. (2018), as well as elastic corrections for free and fixed surface boundary effects. We make the assumption that internal dislocation sources are the residue of the single-crystal growth process. A series of simulations are run at 300K and 900K, with randomly seeded initial dislocation sources including Frank-Read sources, single-arm sources, and mixed sources. At 300K, screw dislocation mobility is limited and many sources are activated to accommodate the imposed strain rate. Plastic flow is maintained by slow and steady propagation of long straight screw dislocations, with periodic emission of fast edge components from the source activation process, consistent with previous simulations Cui et al. (2016); Alcalá et al. (2020), as well as TEM observations Pozuelo et al. (2021). Representative stress-strain curves and PSD of stress fluctuations are shown in Fig. 3. Although the flow stress slightly depends on the specific initial microstructure, at 300K the PSD of all simulations scales as ${}1/f^{2}$. The PSD power law exponent agrees well with experiments, confirming that low-temperature plasticity is a simple superimposition of the uncorrelated movement of short-lived edge and long-lived screw dislocations. As expected, our DDD simulations capture the temperature dependence of the flow stress and the emergence of a mixed and curved dislocation microstructure with increasing temperature Po et al. (2016). However, PSD $P_{\sigma}(f)$ scales as $1/f^{2}$ at all temperatures and no large stress drops are observed in the simulated stress-strain curves, in disagreement with the experimental data. We conclude that another thermally activated process, not included in our simulations, is responsible for the large stress fluctuations observed in experiments at high temperatures.
To identify the nature of this process, we consider the statistics of the number of dislocations that exit the pillars during an avalanche event. The number of dislocation loops that leave a pillar within a stress drop $\Delta\sigma$ is found from the relation $\Delta\sigma=E(\Delta\varepsilon-nb/h)$, where $\Delta\varepsilon$ is the applied strain increment, $\Delta\varepsilon^{p}=nb/h$ is the plastic strain increment, $b$ is the magnitude of the Burgers vector along the loading direction and $h$ is the height of the pillar. Assuming that these $n$ dislocations are emitted near the peak stress of the event, we can construct the CDF of the number of dislocations emitted as a function of stress and temperature, as shown in Fig. 4. A model of this CDF can be developed as follows. Assume that dislocations are emitted at a rate $\nu(\sigma,T)=N\nu_{0}\exp\left(-\Delta H(\sigma)/\,k_{B}T\right)$, where $\nu_{0}$ is the attempt frequency, and $N$ is the number of source sites Zhu et al. (2008). The activation enthalpy $\Delta H(\sigma)=\Delta U_{0}-m\sigma V$ depends on the shear stress $\tau=m\sigma$, where $m$ is the Schmid factor, $\Delta U_{0}$ is the activation energy, and $V$ is the activation volume. Following Chen et al. (2015), we derive the cumulative distribution of dislocations emitted from all sources as $\text{CDF}(\sigma,T)=1-\text{exp}\left(-\frac{1}{E\dot{\varepsilon}}\int_{0}^{\sigma}\nu(\xi,T)\,\text{d}\xi\right)$, which assumes elastic loading between burst events. Using nonlinear least squares regression, we fit this function to the high-temperature experimental data in Fig. 4, and find that $N\nu_{0}\sim 3.6~{}$\mathrm{s}^{-1}$$, $V\sim 8b^{3}$, and $\Delta U_{0}\sim 0.68$ eV. (Fig. 4).
This activation energy is lower than the value expected for the glide of screw dislocations Gumbsch et al. (2002), indicating that mobility is not the only mechanism controlling the transition between uncorrelated and correlated plastic activity in the pillars. Indeed both activation energy and volumes are consistent with a nucleation process, as demonstrated by other experiments and molecular dynamics simulations Zhu et al. (2008); Brochard et al. (2010); Chen et al. (2015). Nucleation is expected to take place from areas of stress concentration on the pillar surface, such as indenter-pillar contact points and FIB-induced nanometer-scale ledges. This was already proposed by Schneider et al., based on the effects of FIB re-milling in Mo micropillars of comparable size compressed at room temperature Schneider et al. (2010). The synergies effects of sufficient dislocation mobility and increased nucleation rate also explain the high temporal and spatial correlation features of avalanche events observed at high temperatures. This is because, given enough mobility, nucleation can act to relieve local stress concentrations by emission of arrays of like dislocations. These coherent dislocations, emitted from a localized source, sweep the crystal on adjacent planes and within a short time window, thus forming localized slip bands and large stress drops. The coincidence of temporal and spatial events is observed in our results, where slip bands become evident after large stress drops (e.g. inset in Fig. 1). Because a proper flow stress is never achieved before a large nucleation event takes place, at temperatures above $600$K this process produces saw-tooth stress-strain profiles, as seen in Fig. 1. Individual source exhaustion caused by local stress relaxation typically leads to the activation of multiple nucleation sources and to the formation of the corresponding slip bands during a single test.
Figure 4: Cumulative probability function of nucleation stress. The dashed lines are model predictions from the supplementary equation 7.
Our results indicate that low-temperature avalanches are the manifestation of uncorrelated plastic activity dominated by low-mobility screw dislocations, showing no indication of critical behavior. A transition to a regime of highly correlated events is observed with increasing temperature. This transition occurs because, together with increased mobility, dislocation nucleation from surface stress concentrations becomes increasingly active. This results in correlated dislocation activity that exhibits features of critical behavior. Given that different BCC micropillars show similar deformation characteristics as a function of their homologous temperatures Schneider et al. (2009); Abad et al. (2016), it is likely that the origin of this transition is a universal feature in BCC micropillars.
We conclude by arguing that the nature of the emergence of this transition with increasing temperature, that is, the increased and synergistic operation of mobility and nucleation in the presence of local stress concentrations, is the same as in the brittle-to-ductile transition (BDT) observed in all bulk BCC metals. In the case of bulk BDT, cracks act as internal stress concentrators, similar to surface stress concentrators in micropillars. That the nature of the two transitions is the same is also suggested by the similarity of certain key parameters such as the transition temperature and energy. In single crystal W, the temperature of the BDT is $\sim 370-470$ K, while its activation energy, measured in both bulk Gumbsch et al. (1998) and microscale Ast et al. (2018) fracture tests, is in the range $\sim 0.2-0.5$ eV. The relation between the two phenomena, once clearly established, has important practical applications because it allows us to screen the BDT properties from avalanche data extracted from inexpensive, high-throughput, and small-scale tests.
Acknowledgements.
G.P. acknowledges the support of the US DOE through grants number DE-SC0024675 and DE-SC0024401, and the US NSF through CAREER grant number 2340174 with the University of Miami. Y.L. is supported by the National Natural Science Foundation of China under Grant No. 12302279, and the support of the U.S. Department of Energy at UCLA is acknowledged under DE-SC0018410:0008.